OpenStudy (anonymous):

Solve the equation by graphing: x2 – 18x + 81 = 0. A. {-2, 9} B. {9/2, 3} C. {9} D. No solution

Directrix (directrix):

C. {9} x2 – 18x + 81 = 0 (x-9)(x-9) = 0 x = 9 or x = 9 --> double root, point of tangency on x-axis